OpenStudy (anonymous):

Alex wants to put a fence around a square garden. The garden has and area of 784 ft2. What is the perimeter of the garden?

OpenStudy (anonymous):

Formula for area of square is l^2, where l is the length. Meaning A=l^2 replacing the area and find l, what do you get?

OpenStudy (anonymous):

I got 112

OpenStudy (anonymous):

No. the formula is 784=l^2, is not it?

OpenStudy (lexi724):

Depends on the shape you are talking about... If its a square, perimeter = 4*√(s) where 's' is the area. If its a rectangle, perimeter = 2*(xy/y + xy/x), 'xy' is their area, 'x' and 'y' are sides. I know this is vague, but the only way you could get it... If its a circle, perimeter = 2/r * a where 'a' is the area. for an equilateral triangle, perimeter = 3*(√s * √4/ª√3), where 's' is the area, and a=4.

OpenStudy (anonymous):

making L subject of the formula, what do you get?

OpenStudy (anonymous):

@lisa2005 ?

OpenStudy (anonymous):

lost

OpenStudy (anonymous):

Okay. I am going to give you the steps on how to make l the subject in the equation of area. Follow this:\[A=l ^{2}\]\[\sqrt{A}=l\]. Clear?

OpenStudy (anonymous):

28*4